write the fraction as a mixed number or a whole number. This is what kind of math I am having trouble with could you please help.
1. 20/5
2. 19/4
3. 22/7
4. 39/10
5. 19/10

Answer by jim_thompson5910(35256)

1.

20/5 = (5*4)/5 = 4

=======================================================
2.

19/4 = (16+3)/4 = 16/4 + 3/4 = 4 + 3/4 = 4 3/4 (this is a mixed number)

=======================================================
3.

22/7 = (21+1)/7 = 21/7 + 1/7 = 3 + 1/7 = 3 1/7 (this is a mixed number)
=======================================================
4.

39/10 = (30+9)/10 = 30/10 + 9/10 = 3 + 9/10 = 3 9/10 (this is a mixed number)
=======================================================
5.

19/10 = (10+9)/10 = 10/10 + 9/10 = 1 + 9/10 = 1 9/10 (this is a mixed number)

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim